### v2

# SemViQA: A Semantic Question Answering System for Vietnamese Information Fact-Checking

Abstract
The rise of misinformation, exacerbated by Large Language Models (LLMs) like GPT and Gemini, demands robust fact-checking solutions, especially for low-resource languages like Vietnamese. Existing methods struggle with semantic ambiguity, homonyms, and complex linguistic structures, often trading accuracy for efficiency. We introduce SemViQA, a novel Vietnamese fact-checking framework integrating Semantic-based Evidence Retrieval (SER) and Two-step Verdict Classification (TVC). Our approach balances precision and speed, achieving state-of-the-art results with 78.97% strict accuracy on ISE-DSC01 and 80.82% on ViWikiFC, securing 1st place in the UIT Data Science Challenge. Additionally, SemViQA Faster improves inference speed 7× while maintaining competitive accuracy. SemViQA sets a new benchmark for Vietnamese fact verification, advancing the fight against misinformation. The source code is available at: https://github.com/DAVID-NGUYEN-S16/SemViQA.
SemViQA: A Semantic Question Answering System for Vietnamese Information Fact-Checking
1 Introduction
The rapid advancement of large language models (LLMs), such as OpenAI’s ChatGPT, Google Gemini Team et al. (2024), Llama3.1 Touvron et al. (2023), Qwen2.5 Qwen et al. (2025), DeepSeek V3, DeepSeek-AI et al. (2024), Phi3.5 Abdin et al. (2024) has significantly improved information retrieval and processing across various domains. However, a major challenge with these systems is their tendency to generate factually incorrect or hallucinated content seemingly plausible information that lacks factual grounding Soleimani et al. (2020). This issue is particularly critical in domains requiring high accuracy, such as healthcare, law, and journalism, where misinformation can have serious consequences. Consequently, developing reliable fact-checking systems capable of retrieving and evaluating evidence from real-world sources has become an urgent need in Natural Language Processing (NLP).
Although fact verification has been extensively studied in high-resource languages like English, with a primary focus on datasets such as FEVER Thorne et al. (2018), LIAR Wang (2017), PubHealthTab Akhtar et al. (2022), and TabFact Chen et al. (2020), extending these methods to low-resource languages like Vietnamese remains a significant challenge. Transformer-based approaches, including BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019), have shown strong performance in fact verification, yet their adaptation to Vietnamese is still underexplored. For Vietnamese fact verification, ViNSV Tran et al. (2024) employs BM25 and SBERT Reimers and Gurevych (2019a) for evidence retrieval. However, SBERT’s 256-token input limitation leads to information loss in long-context claims, making it particularly ineffective for complex cases requiring multi-sentence reasoning or cross-referencing distant evidence. As a result, the system struggles with hard samples that demand deeper contextual understanding, ultimately reducing fact verification accuracy. In contrast, graph-based reasoning methods in English Zhong et al. (2020) have shown notable potential for semantic inference but are often hindered by high computational costs, restricting their real-world applicability. Moreover, traditional retrieval techniques like TF-IDF and BM25, though widely used, rely heavily on token frequency and exact keyword matching, making them less effective at capturing the nuanced semantics of complex claims. While some recent work employing large language models (LLMs) (Huo et al., 2023; Schimanski et al., 2024) has shown promise in addressing hard retrieval questions, these approaches typically require considerably more computing resources and longer processing times compared to conventional methods. Thus, a trade-off persists: models offering high accuracy tend to operate slowly, whereas faster models often compromise on accuracy.
To address the challenges in fact verification, we introduce SemViQA, a system designed to optimize both accuracy and efficiency. It comes in two variants: SemViQA Standard, which prioritizes accuracy, and SemViQA Faster, which significantly improves inference speed while maintaining strong performance. The system is built upon three core components:
1.
Data Processing: A specialized preprocessing pipeline designed to efficiently handle long-token sequences, a known limitation of Transformer-based models. Our approach preserves the semantic integrity of input claims while ensuring optimal tokenization for downstream processing.
2.
Semantic-based Evidence Retrieval (SER): A hybrid retrieval framework combining TF-IDF for efficient keyword-based matching and a Question Answering Token Classifier (QATC) for semantic-based evidence selection. This approach balances speed and accuracy by prioritizing TF-IDF for common cases while activating QATC selectively for hard samples, significantly improving precision while maintaining computational efficiency.
3.
Two-step Verdict Classification (TVC): A hierarchical classification strategy leveraging Focal Loss Lin et al. (2018) and Cross-Entropy Loss to enhance claim verification accuracy. This approach first determines whether a claim is Supported, Refuted, or Not Enough Information (NEI) and subsequently refines the decision boundary for hard-to-classify cases, reducing misclassification and improving robustness.
We evaluate SemViQA on two of the largest Vietnamese fact verification datasets, ISE-DSC01 and ViWikiFC Le et al. (2024). Our system achieves 78.97% strict accuracy on ISE-DSC01 111https://codalab.lisn.upsaclay.fr/competitions/15497#results and 80.82% on ViWikiFC, outperforming existing baselines. These results highlight SemViQA’s potential to improve the accuracy of fact verification in Vietnamese, contributing to misinformation mitigation and enhancing transparency in digital content.
The remainder of this paper is structured as follows: Section 2 discusses related work. Section 3 details our proposed methodology, including SER, TVC, and QATC. Section 4 presents experimental results and performance evaluation. Finally, Section 5 concludes the paper and outlines future research directions.
2 Related Works
Advancements in Natural Language Processing (NLP) have significantly improved information verification and evidence extraction. Early methods, such as the Neural Semantic Matching Network (NSMN) Nie et al. (2018), leveraged BiLSTM Graves and Schmidhuber (2005) and WordNet-based features to enhance verification accuracy but struggled with complex sentence relations due to LSTM’s sequential nature.
The introduction of transformer models, notably BERT Devlin et al. (2019), transformed claim verification by enabling bidirectional contextual encoding Soleimani et al. (2020); Zhou et al. (2019); Malon (2018). While achieving state-of-the-art results on FEVER Aly et al. (2021); Lin et al. (2024); Yuan and Vlachos (2024); DeHaven and Scott (2023), BERT’s input length limitation makes it less effective for long-context fact-checking.
Graph-based models have further advanced multi-step reasoning. Approaches such as GCN-based reasoning Zhong et al. (2020); Thorne et al. (2018) and AdMIRaL’s logic-driven document retrieval Aly and Vlachos (2022) improve evidence sufficiency but demand high computational resources.
For Vietnamese fact verification, research remains limited. ViNSV Tran et al. (2024) combined BM25 and SBERT Reimers and Gurevych (2019a) for evidence retrieval but struggled with complex reasoning due to static embeddings.
Building on prior work, SemViQA optimizes fact verification by integrating fast keyword-based retrieval (TF-IDF) with semantic reasoning (QATC) and a hierarchical classification strategy, ensuring high accuracy, fast inference, and scalability for real-world applications.
3 SemViQA - Semantic Vietnamese Question Answering
Our proposed method, SemViQA, employs a three-stage pipeline to enhance fact verification, as illustrated in Figure 1. The process begins with a preprocessing stage, where input data is restructured to ensure compatibility with retrieval and classification models. Subsequently, evidence retrieval is performed using TF-IDF for straightforward cases, while the Question Answering Token Classifier (QATC) refines evidence selection for more complex claims. Finally, in the claim classification stage, a hierarchical approach integrates Cross-Entropy Loss and Focal Loss Lin et al. (2018) to improve both accuracy and robustness.
3.1 Data Processing
Before implementing data processing solutions, we conducted a thorough analysis of the input data to understand its characteristics and challenges. One of the prominent issues we encountered is the token limit of current Vietnamese BERT models, as clearly illustrated in Figure 2.
During data processing, we observed that the input context has a considerable length, with the longest context extending up to 4803 tokens. This presents a challenge for data processing given the limitations of 256 tokens for classification and 512 tokens for the Question Answering model.
[…]
From a long context, we proceed to split it into smaller segments. We then slide through each segment and check the number of tokens in each segment. If the token count does not exceed 400, we append the segment to the subcontext. When the subcontext has a token count greater than 400 (see Figure 3), we check if the evidence sentence is contained within the subcontext. If it is, then the subcontext is considered to contain that evidence. Otherwise, if it is not present, the evidence for that subcontext is assigned as an empty string. This process allows us to create new data samples with contexts containing fewer than 512 tokens, which is beneficial for training the model more effectively. This ensures that no segment is omitted, unlike stride-based cutting methods, which may result in evidence sentences being split between segments within a complete paragraph.
3.2 Semantic-based Evidence Retrieval (SER)
In addressing this problem, focusing on evidence retrieval is a crucial and decisive part. Accurate classification can only be performed when the evidence is correct. After analyzing and training on the data, we found that methods like TF-IDF and BM25 are effective for simpler samples but struggle with more complex samples.
Therefore, we have divided the evidence retrieval process into two parts:
Part 1: Using TF-IDF: We begin by segmenting the context into smaller passages and pairing each segment with the claim to generate structured data samples, which then undergo a preprocessing pipeline including noise removal and tokenization using ViTokenizer 222https://github.com/trungtv/pyvi to standardize the input. A key observation is that TF-IDF performs well on simple claims, particularly those classified as refuted, but struggles with claims requiring deeper reasoning, as it relies solely on keyword matching without capturing semantic relationships. To address this limitation, we introduce an enhanced strategy to enrich contextual information, where segments containing fewer than 60% of the claim’s token count are merged with the preceding segment to supplement missing information. This is particularly beneficial for short or incomplete sentences where crucial words may be absent from the claim, leading to improved evidence retrieval accuracy. After retrieving potential evidence using TF-IDF, we rank the retrieved segments and apply a Confidence Threshold to classify them into easy and challenging cases, as detailed in Appendix C. While simple cases can be resolved directly using TF-IDF, more complex cases are forwarded to Part 2 for further analysis, ensuring higher classification accuracy.
Part 2: Using the Question Answering Token Classifier Model: The subcontext is also used for model QA inference instead of inferring the full context. In this process, three situations can occur: The first, when multiple subcontexts generate different responses, we interpret this as the model exhibiting confusion in handling evidence responses. In this case, we collect all retrieved evidence and apply TF-IDF to select the most appropriate one. The second, when only one piece of evidence appears across multiple subcontexts, we interpret this as the model being highly confident in its prediction. The third, when the model does not retrieve any evidence, we use TF-IDF to extract relevant evidence as input for our TVC method. By combining these methods, we have achieved the best results in accurately and reliably retrieving evidence.
3.3 Two-step Verdict Classification (TVC)
Our input data consists of claim-evidence pairs, where the claim is the statement that needs to be verified, and the evidence comprises information used to support (supported), refute (refuted), or indicate insufficient information (NEI) regarding the claim. To address this problem, we undertake two main tasks as follows:
3.4 Three-Class Classification
The initial step involves classifying each claim-evidence pair into one of three categories: Supported, Refuted, or Not Enough Information (NEI). We define a three-class classification function as follows:
where denotes the predicted probability distribution over the three classes, and represents a neural network model based on an advanced BERT-derived architecture, fine-tuned for this task. The predicted label is determined by:
where is the class with the highest predicted probability. The model is optimized using Cross-Entropy Loss, ensuring that aligns closely with the ground-truth distribution.
3.5 Binary Classification
When the three-class model does not predict NEI (i.e., ), a secondary binary classification step refines the decision between Supported and Refuted. We define a binary classification function:
where represents the predicted probability distribution over the two classes, and is a neural network based on a fine-tuned BERT-derived architecture, specifically trained for binary classification. The predicted label is given by:
where corresponds to the class with the highest probability. This model is optimized using Focal Loss Lin et al. (2018) to address class imbalance and enhance performance on difficult examples.
3.6 Final Decision Rule
The final predicted class, denoted , is derived by integrating the outputs of the three-class and binary classification models. When the three-class model predicts , the decision is immediate. In all other cases, a comparison is made between the maximum probability from the three-class model (excluding the NEI class) and that from the binary model. This process is formalized by the following decision rule:
where represents the highest probability assigned by the three-class model to either the Supported or Refuted class, and denotes the corresponding maximum probability from the binary model.
This approach leverages the complementary strengths of both models: the three-class model provides a broad categorization, including the critical NEI class, while the binary model offers refined discrimination between Supported and Refuted claims. By prioritizing the three-class prediction when its confidence exceeds that of the binary model, the method ensures robust handling of complex claims while maintaining high accuracy for well-defined cases.
3.7 Question Answering Token Classifier (QATC)
Typical QA models are trained only to predict the start and end positions of an answer. Here, we extend this by introducing a Token Classification task, where we assign a label of 1 to tokens within the answer and 0 otherwise. This enhances the model’s ability to identify relevant tokens for answer extraction.
Procedure: We input the context and claim into the model, extract token features, and apply two linear layers to predict answer start and end probabilities. The default QA loss is the Cross Entropy loss:
where is the number of classes, is the true label, and is the softmax probability for the class.
3.7.1 Token Classification Task
Inspired by Rationale Tagging Ju et al. (2019), we predict whether each token is part of the answer (label 1) or not (label 0). For NEI (Not Enough Information) cases, all tokens are labeled 0. To achieve this, we introduce an additional fully connected (FC) layer that computes classification probabilities:
where is the output representation of token from RoBERTa Zhuang et al. (2021); Aly et al. (2021). The loss for this task is the Binary Cross Entropy loss:
3.7.2 Rationale Regularization Loss
To encourage sparsity and continuity in rationale predictions, we introduce a Rationale Regularization Loss , which consists of:
1. Sparsity Loss: Encourages the model to minimize the number of tokens selected as rationales:
where is a mask ensuring that the loss is computed only on non-padding tokens.
2. Continuity Loss: Encourages smoother rationale selection by minimizing sudden jumps between selected tokens:
The total rationale regularization loss is:
3.7.3 Combined Loss Function
The final objective function integrates all components:
where controls the weight of the rationale tagging loss, and , are hyperparameters for regularization.
4 Experiments
4.1 Dataset
ISE-DSC01, used in the UIT Challenge 2023, is designed for AI evaluation in information extraction and verification. It comprises training, public test, and private test sets with diverse claim-context pairs, posing challenges in evidence retrieval and fact-checking.
ViWikiFC Le et al. (2024) was developed to combat misinformation, featuring over 20,000 claims from Wikipedia. Unlike ISE-DSC01, it explicitly includes evidence for the “nei” (Not Enough Information) label, enhancing real-world applicability.
| ISE-DSC01 | ViWikiFC | |
| Train | 37,967 | 16,738 |
| Dev | 4,794 | 2,090 |
| Test | 5,396 | 2,091 |
Table 1 summarizes the datasets. For ViWikiFC, experiments were conducted on all three sets: training, public test, and private test.
| Method | Parameter | ViWikiFC | ISE-DSC01 | Avg Strict Acc | |||||||
| ER | VC | Strict Acc | VC Acc | ER Acc | Time (s) | Strict Acc | VC Acc | ER Acc | Time (s) | ||
| TF-IDF | InfoXLMlarge | 560M | 75.56 | 82.21 | 90.15 | 131 | 73.59 | 78.08 | 76.61 | 378 | 74.58 |
| XLM-Rlarge | 560M | 76.47 | 82.78 | 90.15 | 134 | 75.61 | 80.50 | 78.58 | 366 | 76.04 | |
| Ernie-Mlarge | 560M | 75.56 | 81.83 | 90.15 | 144 | 78.19 | 81.69 | 80.65 | 403 | 76.88 | |
| BM25 | InfoXLMlarge | 560M | 70.44 | 79.01 | 83.50 | 130 | 72.09 | 77.37 | 75.04 | 320 | 71.27 |
| XLM-Rlarge | 560M | 70.97 | 78.91 | 83.50 | 132 | 73.94 | 79.37 | 76.95 | 333 | 72.46 | |
| Ernie-Mlarge | 560M | 70.21 | 78.29 | 83.50 | 141 | 76.58 | 80.76 | 79.02 | 381 | 73.40 | |
| SBert | InfoXLMlarge | 838M | 74.99 | 81.59 | 89.72 | 195 | 71.20 | 76.59 | 74.15 | 915 | 73.10 |
| XLM-Rlarge | 838M | 75.80 | 82.35 | 89.72 | 194 | 72.85 | 78.78 | 75.89 | 835 | 74.33 | |
| Ernie-Mlarge | 838M | 75.13 | 81.44 | 89.72 | 203 | 75.46 | 79.89 | 77.91 | 920 | 75.30 | |
| QA-based approaches | VC | ||||||||||
| ViMRClarge | InfoXLMlarge | 1120M | 77.28 | 81.97 | 92.49 | 3778 | 54.36 | 64.14 | 56.84 | 9798 | 65.82 |
| XLM-Rlarge | 1120M | 78.29 | 82.83 | 92.49 | 3824 | 53.98 | 66.70 | 57.77 | 9809 | 66.14 | |
| Ernie-Mlarge | 1120M | 77.38 | 81.92 | 92.49 | 3785 | 56.62 | 62.19 | 58.91 | 9833 | 67.00 | |
| InfoXLMlarge | InfoXLMlarge | 1120M | 78.14 | 82.07 | 93.45 | 4092 | 53.50 | 63.83 | 56.17 | 10057 | 65.82 |
| XLM-Rlarge | 1120M | 79.20 | 83.07 | 93.45 | 4096 | 53.32 | 66.70 | 57.25 | 10066 | 66.26 | |
| Ernie-Mlarge | 1120M | 78.24 | 82.21 | 93.45 | 4102 | 56.34 | 62.36 | 58.69 | 10078 | 67.29 | |
| LLM | |||||||||||
| Qwen2.5-1.5B-Instruct | 1.5B | 51.03 | 65.18 | 78.96 | 7665 | 59.23 | 66.68 | 65.51 | 19780 | 55.13 | |
| Qwen2.5-3B-Instruct | 3B | 44.38 | 62.31 | 71.35 | 12123 | 60.87 | 66.92 | 66.10 | 31284 | 52.63 | |
| LLM | VC | ||||||||||
| Qwen2.5-1.5B-Instruct | InfoXLMlarge | 2B | 66.14 | 76.47 | 78.96 | 7788 | 64.40 | 68.37 | 66.49 | 19970 | 65.27 |
| XLM-Rlarge | 2B | 67.67 | 78.10 | 78.96 | 7789 | 64.66 | 69.63 | 66.72 | 19976 | 66.17 | |
| Ernie-Mlarge | 2B | 66.52 | 76.52 | 78.96 | 7794 | 65.70 | 68.37 | 67.33 | 20003 | 66.11 | |
| Qwen2.5-3B-Instruct | InfoXLMlarge | 3.5B | 59.88 | 72.50 | 71.35 | 12246 | 65.72 | 69.66 | 67.51 | 31477 | 62.80 |
| XLM-Rlarge | 3.5B | 60.74 | 73.08 | 71.35 | 12246 | 66.12 | 70.44 | 67.83 | 31483 | 63.43 | |
| Ernie-Mlarge | 3.5B | 60.02 | 72.21 | 71.35 | 12251 | 67.48 | 70.77 | 68.75 | 31512 | 63.80 | |
| SER Faster (ours) | TVC (ours) | ||||||||||
| TF-IDF + ViMRClarge | Ernie-Mlarge | 1680M | 79.44 | 82.93 | 94.60 | 410 | 78.32 | 81.91 | 80.26 | 995 | 78.88 |
| TF-IDF + InfoXLMlarge | 1680M | 79.77 | 83.07 | 95.03 | 487 | 78.37 | 81.91 | 80.32 | 925 | 79.07 | |
| SER (ours) | TVC (ours) | ||||||||||
| TF-IDF + ViMRClarge | InfoXLMlarge | 1680M | 80.25 | 83.84 | 94.69 | 2731 | 75.13 | 79.54 | 76.87 | 5191 | 77.69 |
| XLM-Rlarge | 1680M | 80.34 | 83.64 | 94.69 | 2733 | 76.71 | 81.65 | 78.91 | 5219 | 78.53 | |
| Ernie-Mlarge | 1680M | 79.53 | 82.97 | 94.69 | 2733 | 78.97 | 82.54 | 80.91 | 5225 | 79.25 | |
| TF-IDF + InfoXLMlarge | InfoXLMlarge | 1680M | 80.68 | 83.98 | 95.31 | 3860 | 75.13 | 79.60 | 76.87 | 5175 | 77.91 |
| XLM-Rlarge | 1680M | 80.82 | 83.88 | 95.31 | 3843 | 76.74 | 81.71 | 78.95 | 5200 | 78.78 | |
| Ernie-Mlarge | 1680M | 80.06 | 83.17 | 95.31 | 3891 | 78.97 | 82.49 | 80.91 | 5297 | 79.52 | |
4.2 Experimental Setup
We optimized our model through extensive experiments on an NVIDIA A100 GPU, fine-tuning key hyperparameters while maintaining consistent settings across all runs. The final hyperparameters, selected via rigorous validation, improved both accuracy and strict accuracy on ISE-DSC01 and ViWikiFC. Details are provided in Appendix B. For evaluation, all methods were tested on an NVIDIA T4 instance in Kaggle to ensure fairness.
Additionally, the large language model was fine-tuned on a distributed setup with A100 GPUs, utilizing a structured prompt reformulation process. The dataset, originally in raw text and labels, was transformed into a prompt-based format optimized for LLM fine-tuning. This ensured effective task-specific learning while aligning with pre-trained knowledge. Further details on training setup, prompt design, and preprocessing are in Appendix B.
4.3 Main Results
The results in Table 2 demonstrate that SemViQA outperforms previous methods in Vietnamese fact-checking tasks. Specifically, our model achieves the highest Strict Accuracy, reaching 80.82% on ViWikiFC and 78.97% on ISE-DSC01, establishing a new benchmark for automated fact-checking systems in Vietnamese language.
4.3.1 Performance Comparison
a) Handling Long Token Sequences in Fact-Checking
One major limitation of conventional Question Answering (QA) models in fact verification tasks is their inability to process long-context claims due to token length constraints. Transformer-based models such as ViMRClarge, InfoXLMlarge Chi et al. (2021), and XLM-Rlarge Conneau et al. (2020), Ernie-Mlarge Ouyang et al. (2021) are restricted to a maximum input length of 512 tokens. However, real-world fact-checking datasets, such as ISE-DSC01, often contain contexts exceeding 4800 tokens, causing a significant drop in performance for QA-based approaches due to their inability to handle long-token sequences. This limitation prevents these models from effectively leveraging the full evidence available, leading to incomplete or inaccurate fact verification.
To address this issue, our proposed SemViQA model incorporates a retrieval-based approach capable of handling long-token sequences efficiently, as illustrated in Figure 3. On ISE-DSC01, where long contexts are essential for verification, SemViQA significantly outperforms traditional QA models by effectively capturing and utilizing extended contextual information. This demonstrates that the long-token issue is a critical bottleneck in fact verification tasks.
In contrast, on the ViWikiFC dataset, where the token length is closer to 512 tokens, QA-based models achieve competitive performance, further confirming that the long-token issue is a major factor in performance degradation. However, when the long-token limitation does not exist on the ViWikiFC dataset, integrating our SER approach into a QA-based model proves to be an effective solution. By leveraging SER, we achieve approximately 1.86% higher ER Accuracy on ViWikiFC, demonstrating its efficiency in optimizing fact verification when token constraints are minimized.
These results highlight that long-token constraints are a fundamental challenge in fact verification. Our approach not only mitigates this issue through retrieval-based processing in SemViQA but also enhances QA models with QACT when long-token constraints are less severe, ensuring improved performance across diverse datasets.
b) Performance and Inference Time Optimization
One of the key advancements of SemViQA over other methods is its significant reduction in inference time without compromising accuracy. Key observations include:
The average inference time of SemViQA on ISE-DSC01 is 5200s, whereas large LLM-based models such as Qwen2.5-3B-Instruct Qwen et al. (2024) require over 31,000s, making SemViQA at least 6 times faster than these large language model approaches.
Compared to ViMRClarge333https://huggingface.co/nguyenvulebinh/vi-mrc-large (9800s on ISE-DSC01), SemViQA reduces inference time by nearly 50% while maintaining higher performance in both Strict Accuracy and Veracity Classification Accuracy.
Methods based on SBERT Reimers and Gurevych (2019b) or BM25 have lower inference times but fail to maintain high accuracy, particularly when dealing with multi-step reasoning tasks. SemViQA achieves a better balance between classification performance and processing speed, ensuring practical deployment in real-world environments.
SemViQA Faster: We introduce an optimized version of SemViQA that achieves up to 7× faster inference while maintaining high accuracy. As shown in Figure 4, the inference speed of SER-Faster closely matches that of traditional methods while still outperforming competing solutions in both efficiency and accuracy. To achieve this speed improvement, we process all subcontexts within a sample as a batch, enabling batch inference (see Section 3.1 for details on subcontexts). While this approach results in a slight performance drop compared to the standard SemViQA model, the trade-off is minimal considering the significant reduction in processing time. This efficiency enhancement makes SemViQA Faster highly suitable for real-world deployment, allowing seamless integration into large-scale fact-checking applications.
4.3.2 Analysis of Confidence Threshold in SemViQA
The confidence threshold plays a crucial role in balancing accuracy and inference time in SemViQA’s evidence retrieval process. Analysis from Figure 5 indicates that as the threshold increases from 0.0 to 0.5, evidence retrieval accuracy improves significantly, particularly on ViWikiFC ( 95%) and ISE-DSC01 ( 80.8%). However, beyond 0.5, accuracy gains plateau, while inference time decreases sharply due to the system filtering out low-confidence evidence more aggressively. Setting an optimal threshold in the range of 0.4 – 0.5 achieves a trade-off between efficiency and accuracy, ensuring that SemViQA operates swiftly while maintaining precise evidence retrieval.
4.3.3 Comparison with other results in the competition
| Methods | Strict Acc | VC Acc | ER Acc |
| SemViQA | 78.97 | 82.54 | 80.91 |
| DS@UIT Dynasty | 78.05 | 84.76 | 80.13 |
| URA_FNU | 77.87 | 83.71 | 79.96 |
| Plain Sailing | 77.09 | 82.33 | 78.31 |
| ViNSV | 76.33 | 81.67 | 78.11 |
The results presented in Table 3 indicate that our SemViQA approach outperforms other competing teams, achieving the highest Strict Accuracy and demonstrating exceptional effectiveness in information processing and verification. This achievement highlights SemViQA’s capability to deliver significantly more accurate and reliable results.
5 Conclusion and Future Works
We introduced SemViQA, a Vietnamese fact-checking framework that integrates Semantic-based Evidence Retrieval (SER) and Two-step Verdict Classification (TVC) to enhance claim verification. Our approach outperforms existing methods, including LLMs, TF-IDF, BM25, SBERT, and QA-based models, particularly in handling long-token sequences and complex reasoning tasks. Extensive experiments demonstrated SemViQA’s state-of-the-art performance on ISE-DSC01 and ViWikiFC. Additionally, the SemViQA Faster variant accelerates inference by up to 7×, improving its practicality for real-world applications. By addressing key challenges such as semantic ambiguity and multi-step reasoning, SemViQA lays the groundwork for advancing Vietnamese NLP, with potential applications in misinformation detection and low-resource language fact-checking.
Limitations
While SemViQA demonstrates strong performance in Vietnamese fact verification, several limitations remain. First, our reliance on TF-IDF for initial evidence retrieval, while efficient, limits the model’s ability to capture deep semantic relationships and retrieve implicit evidence. To mitigate this, we employ a threshold-based mechanism to identify hard samples and process them with a more advanced retrieval model. However, this approach relies on manually defined thresholds, which may not generalize well across different datasets, underscoring the need for adaptive and data-driven retrieval strategies in future work. Second, our Two-step Verdict Classification (TVC) framework improves claim verification accuracy but requires multiple classification stages, increasing inference time compared to single-step approaches. This additional computational cost is particularly significant in three-class classification tasks, where optimizing model efficiency without compromising accuracy remains a key challenge. Future work should focus on refining retrieval mechanisms and classification strategies to enhance efficiency and robustness, ensuring broader applicability of SemViQA in real-world fact verification scenarios.
Appendix A Strict Accuracy in Fact-Checking
Strict Accuracy: This metric is a stringent measure that requires both the verdict and the evidence to be predicted correctly compared to the ground truth sample.
Verdict ( and ): refers to the verdict of the sample and the predicted verdict (supported, refuted, nei).
Evidence ( and ): refers to the evidence of the sample and the predicted evidence.
Where:
Strict accuracy is the average of all StrAcc values.
Appendix B Hyperparameter and LLM Training Configuration
In this section, we present the detailed hyperparameter settings and training configurations for both our SemViQA models and the Large Language Model (LLM) fine-tuning process. Table 4 consolidates all hyperparameters used across different models, including Binary Classification (BC), Three-Class Classification (TC), Question Answering Token Classifier (QATC), and LLM fine-tuning.
| Hyperparameter | BC | TC | QATC | LLM |
| Epochs | 20 | 20 | 20 | 1 |
| RT Loss | - | - | ✓ | - |
| Cross-Entropy Loss | - | ✓ | ✓ | - |
| Focal Loss | ✓ | - | - | - |
| Learning Rate | ||||
| Batch Size | 104 | 104 | 36 | 2 |
| Gradient Accumulation | 1 | 1 | 2 | 1 |
| Optimizer (AdamW) | ✓ | ✓ | ✓ | ✓ |
| Max Token Length | 256 | 256 | 512 | 4096 |
| GPUs | A100 | A100 | A100 | A100 |
| Zero | - | - | - | Zero3 |
| LR Schedule | Linear | Linear | Cyclic | Cosine |
| Mixed Precision | - | - | - | bf16 |
We fine-tune a Large Language Model (LLM) using a restructured version of the original datasets, ViWikiFC and ISE-DSC01, as detailed in Figure 5. These datasets have been carefully adapted for training to improve performance and ensure compatibility with our model. For training, we utilize the official Qwen LLM implementation from the QwenLM repository444https://github.com/QwenLM/Qwen. Our training setup follows the full configuration outlined in Table 4, ensuring optimal efficiency and alignment with best practices.
[backgroundcolor=gray!5, linewidth=0.75pt, roundcorner=5pt]
Question: You are tasked with verifying the correctness of the following statement.
- We provide you with a claim and a context. Please classify the claim into one of three labels: “SUPPORTED”, “REFUTED”, or “NEI” (Not Enough Info).
- Your answer should include the classification label and the most relevant evidence sentence from the context.
- Remember, the evidence must be a full sentence, not part of a sentence or less than one sentence.
Given a claim and context as follows:
Context: The actress revealed her secrets to maintaining a youthful appearance as follows: Eating three balanced meals a day. For dinner, Ivy Chen usually eats early to ensure her body has enough time to digest food, metabolize energy, and avoid putting pressure on the stomach and other organs. A recent study published in *Frontiers in Nutrition* suggests that eating dinner earlier can lead to a longer lifespan, with the ideal time being 7 PM. If this is not possible, experts recommend having the last meal of the day 2-3 hours before bedtime. Drinking ginger tea: To keep her body warm, promote blood circulation, and enhance circulation, Ivy Chen drinks ginger tea daily. Her ginger tea is typically made with ground ginger, black tea, turmeric powder, and brown sugar. This drink is a natural remedy that not only boosts the immune system and reduces inflammation but also fights oxidation, supports weight loss, improves skin health, and helps maintain a youthful look. Regular exercise: Ivy Chen is a fitness enthusiast who loves physical activities and exercises daily, even during pregnancy. The Taiwanese actress shared that if she is not busy with work, she runs for at least 30 minutes every day. Even when traveling abroad, she maintains her running habit. A recent study published in *Progress in Cardiovascular Disease* found that regular runners live three years longer than non-runners. Running significantly helps with weight loss, maintaining a balanced physique, toning muscles, relaxing the mind, and benefiting heart health. Besides running, Ivy Chen also swims, practices yoga, and hikes to maintain physical fitness and endurance. Skincare: Regarding her skincare routine, the actress emphasized the importance of hydration. The Taiwanese beauty revealed that she always carries a facial mist to ensure her skin stays hydrated while outdoors.
Claim: Even when traveling abroad, Ivy Chen maintains her running habit.
Answer: This claim is classified as SUPPORTED. The evidence is: Even when traveling abroad, she maintains her running habit.
We present the complete training progress of the LLM models and QATC in Figure 7 and Figure 6, respectively. Figure 7 illustrates the training dynamics of Qwen 1.5B and Qwen 3B, supporting the results presented in Table 2. Notably, the Qwen 1.5B model demonstrates more stable training dynamics compared to the Qwen 3B model during the initial stage. Meanwhile, Figure 6 showcases the completion of QATC training, depicting the loss curves of ViMRC and InfoXLM. These results highlight the convergence behavior of QATC training across different architectures, further supporting the robustness of our approach.
Appendix C Comparison of TF-IDF and QATC in Fact-Checking: Examples of Incorrect vs. Correct Evidence Selection
| Claim | Evidence | TF-IDF | QATC |
Du lịch Triều Tiên là điều mà chỉ có một số người được đi đến.
(Traveling to North Korea is something only a few people can do.) |
Theo nguyên tắc, bất kỳ ai cũng được phép du lịch tới Triều Tiên, và những ai có thể hoàn thành quá trình làm thủ tục thì đều không bị Triều Tiên từ chối cho nhập cảnh.
(In principle, anyone is allowed to travel to North Korea, and those who complete the process are not denied entry.) |
Khách du lịch không được đi thăm thú bên ngoài vùng đã được cho phép trước mà không được hướng dẫn viên người Triều Tiên cho phép nhằm tránh các điệp viên nằm vùng.
(Tourists are not allowed to visit areas outside of the designated zones without a North Korean guide to prevent undercover spies.) |
Theo nguyên tắc, bất kỳ ai cũng được phép du lịch tới Triều Tiên, và những ai có thể hoàn thành quá trình làm thủ tục thì đều không bị Triều Tiên từ chối cho nhập cảnh.
(In principle, anyone is allowed to travel to North Korea, and those who complete the process are not denied entry.) |
Nó có độ nóng chảy ở mức gần 30 độ C.
(It has a melting point of about 30°C.) |
Nó là một kim loại kiềm mềm, màu bạc, và với điểm nóng chảy là 28 °C (83 °F) khiến cho nó trở thành một trong các kim loại ở dạng lỏng tại hay gần nhiệt độ phòng.
(It is a soft, silvery alkali metal with a melting point of 28°C (83°F), making it one of the metals that is liquid at or near room temperature.) |
Nó là nguyên tố có độ âm điện thấp thứ hai sau franci, và chỉ có một đồng vị bền là caesi-133.
(It is the second least electronegative element after francium, and has only one stable isotope, cesium-133.) |
Nó là một kim loại kiềm mềm, màu bạc, và với điểm nóng chảy là 28 °C (83 °F) khiến cho nó trở thành một trong các kim loại ở dạng lỏng tại hay gần nhiệt độ phòng.
(It is a soft, silvery alkali metal with a melting point of 28°C (83°F), making it one of the metals that is liquid at or near room temperature.) |

### v3

@@
 Abstract
-The rise of misinformation, exacerbated by Large Language Models (LLMs) like GPT and Gemini, demands robust fact-checking solutions, especially for low-resource languages like Vietnamese.
+Recent advances in LLMs have accelerated both information generation and misinformation, especially in low-resource languages like Vietnamese, motivating robust fact-checking systems.
 Existing methods struggle with semantic ambiguity, homonyms, and complex linguistic structures, often trading accuracy for efficiency.
@@
 Consequently, developing reliable fact-checking systems capable of retrieving and evaluating evidence from real-world sources has become an urgent need in Natural Language Processing (NLP).
-Although fact verification has been extensively studied in high-resource languages like English, with a primary focus on datasets such as FEVER Thorne et al.
-(2018), LIAR Wang (2017), PubHealthTab Akhtar et al.
-(2022), and TabFact Chen et al.
-(2020), extending these methods to low-resource languages like Vietnamese remains a significant challenge.
-Transformer-based approaches, including BERT Devlin et al.
+Although fact verification has been widely studied in high-resource languages like English, applying these methods to low-resource languages such as Vietnamese remains a significant challenge.
+Transformer-based models, including BERT Devlin et al.
 (2019) and RoBERTa Liu et al.
-(2019), have shown strong performance in fact verification, yet their adaptation to Vietnamese is still underexplored.
-For Vietnamese fact verification, ViNSV Tran et al.
-(2024) employs BM25 and SBERT Reimers and Gurevych (2019a) for evidence retrieval.
-However, SBERT’s 256-token input limitation leads to information loss in long-context claims, making it particularly ineffective for complex cases requiring multi-sentence reasoning or cross-referencing distant evidence.
-As a result, the system struggles with hard samples that demand deeper contextual understanding, ultimately reducing fact verification accuracy.
-In contrast, graph-based reasoning methods in English Zhong et al.
-(2020) have shown notable potential for semantic inference but are often hindered by high computational costs, restricting their real-world applicability.
-Moreover, traditional retrieval techniques like TF-IDF and BM25, though widely used, rely heavily on token frequency and exact keyword matching, making them less effective at capturing the nuanced semantics of complex claims.
-While some recent work employing large language models (LLMs) (Huo et al., 2023; Schimanski et al., 2024) has shown promise in addressing hard retrieval questions, these approaches typically require considerably more computing resources and longer processing times compared to conventional methods.
-Thus, a trade-off persists: models offering high accuracy tend to operate slowly, whereas faster models often compromise on accuracy.
-To address the challenges in fact verification, we introduce SemViQA, a system designed to optimize both accuracy and efficiency.
-It comes in two variants: SemViQA Standard, which prioritizes accuracy, and SemViQA Faster, which significantly improves inference speed while maintaining strong performance.
-The system is built upon three core components:
+(2019), have demonstrated strong performance but their adaptation to Vietnamese is still limited.
+ViNSV Tran et al.
+(2024b) employs BM25 and SBERT Reimers and Gurevych (2019a) for evidence retrieval but suffers from SBERT’s 256-token input constraint, making it ineffective for complex, long-context claims.
+Graph-based reasoning methods Zhong et al.
+(2020) offer promising semantic inference but are often computationally expensive.
+Traditional retrieval methods like TF-IDF and BM25, while efficient, rely heavily on exact keyword matching, limiting their ability to capture nuanced semantics.
+Recent large language model (LLM) approaches Huo et al.
+(2023); Schimanski et al.
+(2024) show potential but typically require substantial computational resources, creating a trade-off between speed and accuracy.
+To address these challenges, we propose SemViQA, a Vietnamese fact-checking framework that balances semantic accuracy and computational efficiency.
+As shown in Figure 1, SemViQA comprises three key components:
 1.
-Data Processing: A specialized preprocessing pipeline designed to efficiently handle long-token sequences, a known limitation of Transformer-based models.
-Our approach preserves the semantic integrity of input claims while ensuring optimal tokenization for downstream processing.
+Semantic-based Evidence Retrieval (SER): Includes a preprocessing step that efficiently handles long-token contexts by splitting them into manageable subcontexts (e.g., 400 tokens).
+It combines fast TF-IDF retrieval with selective Question Answering Token Classification (QATC) to strike a balance between speed and semantic accuracy.
 2.
-Semantic-based Evidence Retrieval (SER): A hybrid retrieval framework combining TF-IDF for efficient keyword-based matching and a Question Answering Token Classifier (QATC) for semantic-based evidence selection.
-This approach balances speed and accuracy by prioritizing TF-IDF for common cases while activating QATC selectively for hard samples, significantly improving precision while maintaining computational efficiency.
-3.
-Two-step Verdict Classification (TVC): A hierarchical classification strategy leveraging Focal Loss Lin et al.
-(2018) and Cross-Entropy Loss to enhance claim verification accuracy.
-This approach first determines whether a claim is Supported, Refuted, or Not Enough Information (NEI) and subsequently refines the decision boundary for hard-to-classify cases, reducing misclassification and improving robustness.
-We evaluate SemViQA on two of the largest Vietnamese fact verification datasets, ISE-DSC01 and ViWikiFC Le et al.
+Two-step Verdict Classification (TVC): Employs a hierarchical classification strategy with both three-class and binary classification stages to enhance robustness and improve performance on challenging claim verification tasks.
+SemViQA achieves 78.97% strict accuracy on ISE-DSC01 111https://codalab.lisn.upsaclay.fr/competitions/15497#results and 80.82% on ViWikiFC Le et al.
+(2024), outperforming existing baselines (see Table 1).
+These results validate SemViQA’s potential to enhance Vietnamese fact verification, supporting misinformation mitigation and improved transparency.
+The rest of this paper is organized as follows: Section 2 reviews related work, Section 3 presents the methodology, Section 4 reports experimental results, and Section 5 concludes the paper with future directions.
+2 Related Works
+Advances in Natural Language Processing (NLP) have driven rapid progress in fact verification and evidence extraction.
+Early BiLSTM‑based models such as the Neural Semantic Matching Network (NSMN) Nie et al.
+(2018)—augmented with WordNet features improved accuracy but struggled with complex sentence relations due to sequential limitations Graves and Schmidhuber (2005).
+Transformer models, notably BERT Devlin et al.
+(2019), introduced bidirectional contextual encoding and achieved state‑of‑the‑art results on FEVER Soleimani et al.
+(2020); Zhou et al.
+(2019); Malon (2018); Aly et al.
+(2021); Lin et al.
+(2024); Yuan and Vlachos (2024); DeHaven and Scott (2023), yet their input‑length cap hinders long‑document fact‑checking.
+Graph‑based reasoning further enhances multi‑hop verification GCN variants Zhong et al.
+(2020); Thorne et al.
+(2018) and AdMIRaL’s logic‑driven retrieval Aly and Vlachos (2022) boost evidence sufficiency but at considerable computational cost.
+Vietnamese research is sparse: ViNSV Tran et al.
+(2024b) pairs BM25 with SBERT Reimers and Gurevych (2019a) yet falters on complex reasoning due to static embeddings.
+TF-IDF remains the industry standard for document retrieval due to its speed, simplicity, interpretability, and ability to handle long contexts effectively, making it suitable for rapid and scalable retrieval Reddy et al.
+(2018); Qaiser and Ali (2018); Li (2021); Azevedo et al.
 (2022).
-Our system achieves 78.97% strict accuracy on ISE-DSC01 111https://codalab.lisn.upsaclay.fr/competitions/15497#results and 80.82% on ViWikiFC, outperforming existing baselines.
-These results highlight SemViQA’s potential to improve the accuracy of fact verification in Vietnamese, contributing to misinformation mitigation and enhancing transparency in digital content.
-The remainder of this paper is structured as follows: Section 2 discusses related work.
-Section 3 details our proposed methodology, including SER, TVC, and QATC.
-Section 4 presents experimental results and performance evaluation.
-Finally, Section 5 concludes the paper and outlines future research directions.
-2 Related Works
-Advancements in Natural Language Processing (NLP) have significantly improved information verification and evidence extraction.
-Early methods, such as the Neural Semantic Matching Network (NSMN) Nie et al.
-(2018), leveraged BiLSTM Graves and Schmidhuber (2005) and WordNet-based features to enhance verification accuracy but struggled with complex sentence relations due to LSTM’s sequential nature.
-The introduction of transformer models, notably BERT Devlin et al.
-(2019), transformed claim verification by enabling bidirectional contextual encoding Soleimani et al.
-(2020); Zhou et al.
-(2019); Malon (2018).
-While achieving state-of-the-art results on FEVER Aly et al.
-(2021); Lin et al.
-(2024); Yuan and Vlachos (2024); DeHaven and Scott (2023), BERT’s input length limitation makes it less effective for long-context fact-checking.
-Graph-based models have further advanced multi-step reasoning.
-Approaches such as GCN-based reasoning Zhong et al.
-(2020); Thorne et al.
-(2018) and AdMIRaL’s logic-driven document retrieval Aly and Vlachos (2022) improve evidence sufficiency but demand high computational resources.
-For Vietnamese fact verification, research remains limited.
-ViNSV Tran et al.
-(2024) combined BM25 and SBERT Reimers and Gurevych (2019a) for evidence retrieval but struggled with complex reasoning due to static embeddings.
-Building on prior work, SemViQA optimizes fact verification by integrating fast keyword-based retrieval (TF-IDF) with semantic reasoning (QATC) and a hierarchical classification strategy, ensuring high accuracy, fast inference, and scalability for real-world applications.
+However, its reliance on surface-level keyword matching limits its capability to handle paraphrases, contextual nuances, and multi-hop reasoning, reducing recall accuracy on complex queries.
+Ensemble learning Hannichenko et al.
+(2023); Wang et al.
+(2021); Liu et al.
+(2024); Ganaie et al.
+(2022) mitigates individual model weaknesses by aggregating diverse architectures and training signals, yielding robust gains.
+Building on these insights, SemViQA fuses fast TF‑IDF retrieval, semantic reasoning via QATC, and hierarchical classification, delivering high accuracy, low latency, and practical scalability for Vietnamese fact verification.
 3 SemViQA - Semantic Vietnamese Question Answering
-Our proposed method, SemViQA, employs a three-stage pipeline to enhance fact verification, as illustrated in Figure 1.
-The process begins with a preprocessing stage, where input data is restructured to ensure compatibility with retrieval and classification models.
-Subsequently, evidence retrieval is performed using TF-IDF for straightforward cases, while the Question Answering Token Classifier (QATC) refines evidence selection for more complex claims.
-Finally, in the claim classification stage, a hierarchical approach integrates Cross-Entropy Loss and Focal Loss Lin et al.
-(2018) to improve both accuracy and robustness.
-3.1 Data Processing
-Before implementing data processing solutions, we conducted a thorough analysis of the input data to understand its characteristics and challenges.
-One of the prominent issues we encountered is the token limit of current Vietnamese BERT models, as clearly illustrated in Figure 2.
-During data processing, we observed that the input context has a considerable length, with the longest context extending up to 4803 tokens.
-This presents a challenge for data processing given the limitations of 256 tokens for classification and 512 tokens for the Question Answering model.
-[…]
-[…]
-From a long context, we proceed to split it into smaller segments.
-We then slide through each segment and check the number of tokens in each segment.
-If the token count does not exceed 400, we append the segment to the subcontext.
-When the subcontext has a token count greater than 400 (see Figure 3), we check if the evidence sentence is contained within the subcontext.
-If it is, then the subcontext is considered to contain that evidence.
-Otherwise, if it is not present, the evidence for that subcontext is assigned as an empty string.
-This process allows us to create new data samples with contexts containing fewer than 512 tokens, which is beneficial for training the model more effectively.
-This ensures that no segment is omitted, unlike stride-based cutting methods, which may result in evidence sentences being split between segments within a complete paragraph.
+We formulate Vietnamese fact verification as a multi-output classification task, where the input is a pair , with being a claim and its corresponding context paragraph or document.
+The objective is to (i) identify the most relevant evidence sentence(s) from and (ii) predict the veracity label of the claim as one of three categories: Supported, Refuted, or Not Enough Information (NEI).
+To address challenges such as long input sequences and semantic ambiguity, we propose SemViQA, a three-stage framework consisting of data pre-processing, semantic-based evidence retrieval, and two-step verdict classification.
+An overview of the architecture is shown in Figure 2, with detailed descriptions provided in the following subsections.
+3.1 Evidence Extraction via Question Answering with Token Classification
+3.1.1 Data Processing
+To effectively support evidence retrieval and claim classification, we apply distinct preprocessing strategies to the context based on the specific requirements of each downstream task (see Appendix A).
+A key challenge arises from the considerable length of many context passages, frequently exceeding the token limits of Vietnamese BERT-based models.
+Figure 3 illustrates our approach for handling long input contexts.
+First, the context is segmented into individual sentences.
+Next, sentences are sequentially aggregated into subcontexts until reaching approximately 400 tokens.
+Each completed subcontext is then processed by the QATC model to identify potential evidence.
+The next subcontext begins from the subsequent sentence, and this process continues until all sentences are processed.
+However, processing subcontexts sequentially can be time-consuming.
+Therefore, we developed SemViQA Faster, which batches and processes subcontexts in parallel, significantly accelerating the retrieval process.
+3.1.2 Question Answering with Token Classification (QATC)
+Traditional Question Answering models typically predict the start and end positions of an answer span.
+In our framework, we enhance this approach by introducing a token-level classification objective, enabling the model to focus explicitly on tokens within evidence sentences in the context.
+This dual formulation provides improved supervision for evidence extraction.
+Drawing inspiration from rationale tagging Ju et al.
+(2019), we treat token labeling as a binary classification task: tokens within evidence sentences receive a label of 1, and all other tokens receive a label of 0.
+In cases marked NEI (Not Enough Information), every token is labeled as 0.
+We employ a feed-forward classification layer on the token representations:
+where is the contextual representation of token , are learnable weights in the neural network and denotes the sigmoid function. The loss for this task is the Binary Cross-Entropy (BCE) loss:
 3.2 Semantic-based Evidence Retrieval (SER)
-In addressing this problem, focusing on evidence retrieval is a crucial and decisive part.
-Accurate classification can only be performed when the evidence is correct.
-After analyzing and training on the data, we found that methods like TF-IDF and BM25 are effective for simpler samples but struggle with more complex samples.
-Therefore, we have divided the evidence retrieval process into two parts:
-Part 1: Using TF-IDF: We begin by segmenting the context into smaller passages and pairing each segment with the claim to generate structured data samples, which then undergo a preprocessing pipeline including noise removal and tokenization using ViTokenizer 222https://github.com/trungtv/pyvi to standardize the input.
-A key observation is that TF-IDF performs well on simple claims, particularly those classified as refuted, but struggles with claims requiring deeper reasoning, as it relies solely on keyword matching without capturing semantic relationships.
-To address this limitation, we introduce an enhanced strategy to enrich contextual information, where segments containing fewer than 60% of the claim’s token count are merged with the preceding segment to supplement missing information.
-This is particularly beneficial for short or incomplete sentences where crucial words may be absent from the claim, leading to improved evidence retrieval accuracy.
-After retrieving potential evidence using TF-IDF, we rank the retrieved segments and apply a Confidence Threshold to classify them into easy and challenging cases, as detailed in Appendix C.
-While simple cases can be resolved directly using TF-IDF, more complex cases are forwarded to Part 2 for further analysis, ensuring higher classification accuracy.
-Part 2: Using the Question Answering Token Classifier Model: The subcontext is also used for model QA inference instead of inferring the full context.
-In this process, three situations can occur: The first, when multiple subcontexts generate different responses, we interpret this as the model exhibiting confusion in handling evidence responses.
-In this case, we collect all retrieved evidence and apply TF-IDF to select the most appropriate one.
-The second, when only one piece of evidence appears across multiple subcontexts, we interpret this as the model being highly confident in its prediction.
-The third, when the model does not retrieve any evidence, we use TF-IDF to extract relevant evidence as input for our TVC method.
-By combining these methods, we have achieved the best results in accurately and reliably retrieving evidence.
+Accurate claim verification requires reliable evidence. To improve both efficiency and robustness, we adopt a two-stage evidence retrieval strategy combining TF-IDF with a QATC.
+Stage 1: TF-IDF-based Retrieval.
+We segment the context into smaller passages and pair each with the claim .
+Preprocessing includes noise removal and tokenization using ViTokenizer222https://github.com/trungtv/pyvi.
+TF-IDF is effective for simple claims particularly refuted ones but struggles with semantically complex cases due to its reliance on keyword overlap.
+To enrich short segments (i.e., those with fewer than 60% of ’s tokens), we merge them with preceding segments to improve evidence completeness.
+Retrieved segments are then ranked, and a confidence threshold is applied to identify easy cases (handled by TF-IDF) and hard cases (passed to QATC).
+Stage 2: QATC-based Refinement.
+For complex cases, QATC is applied to segmented subcontexts rather than the full context due to the input length limitation of BERT models.
+The detailed processing approach is described in Section 3.1.1.
+At this time, we consider three scenarios: (1) If multiple subcontexts yield conflicting answers, we collect all predicted spans and re-rank them using TF-IDF.
+(2) If a single evidence span consistently appears, it is selected directly.
+(3) If no evidence is found, fallback to TF-IDF is used.
+This hybrid approach balances speed and semantic accuracy, improving evidence selection for downstream verdict classification. Examples are provided in Appendix E.
 3.3 Two-step Verdict Classification (TVC)
-Our input data consists of claim-evidence pairs, where the claim is the statement that needs to be verified, and the evidence comprises information used to support (supported), refute (refuted), or indicate insufficient information (NEI) regarding the claim.
-To address this problem, we undertake two main tasks as follows:
-3.4 Three-Class Classification
-The initial step involves classifying each claim-evidence pair into one of three categories: Supported, Refuted, or Not Enough Information (NEI).
-We define a three-class classification function as follows:
-where denotes the predicted probability distribution over the three classes, and represents a neural network model based on an advanced BERT-derived architecture, fine-tuned for this task.
-The predicted label is determined by:
-where is the class with the highest predicted probability. The model is optimized using Cross-Entropy Loss, ensuring that aligns closely with the ground-truth distribution.
-3.5 Binary Classification
-When the three-class model does not predict NEI (i.e., ), a secondary binary classification step refines the decision between Supported and Refuted. We define a binary classification function:
-where represents the predicted probability distribution over the two classes, and is a neural network based on a fine-tuned BERT-derived architecture, specifically trained for binary classification.
-The predicted label is given by:
-where corresponds to the class with the highest probability. This model is optimized using Focal Loss Lin et al. (2018) to address class imbalance and enhance performance on difficult examples.
-3.6 Final Decision Rule
-The final predicted class, denoted , is derived by integrating the outputs of the three-class and binary classification models.
-When the three-class model predicts , the decision is immediate.
-In all other cases, a comparison is made between the maximum probability from the three-class model (excluding the NEI class) and that from the binary model.
-This process is formalized by the following decision rule:
-where represents the highest probability assigned by the three-class model to either the Supported or Refuted class, and denotes the corresponding maximum probability from the binary model.
-This approach leverages the complementary strengths of both models: the three-class model provides a broad categorization, including the critical NEI class, while the binary model offers refined discrimination between Supported and Refuted claims.
-By prioritizing the three-class prediction when its confidence exceeds that of the binary model, the method ensures robust handling of complex claims while maintaining high accuracy for well-defined cases.
-3.7 Question Answering Token Classifier (QATC)
-Typical QA models are trained only to predict the start and end positions of an answer.
-Here, we extend this by introducing a Token Classification task, where we assign a label of 1 to tokens within the answer and 0 otherwise.
-This enhances the model’s ability to identify relevant tokens for answer extraction.
-Procedure: We input the context and claim into the model, extract token features, and apply two linear layers to predict answer start and end probabilities.
-The default QA loss is the Cross Entropy loss:
-where is the number of classes, is the true label, and is the softmax probability for the class.
-3.7.1 Token Classification Task
-Inspired by Rationale Tagging Ju et al.
-(2019), we predict whether each token is part of the answer (label 1) or not (label 0).
-For NEI (Not Enough Information) cases, all tokens are labeled 0.
-To achieve this, we introduce an additional fully connected (FC) layer that computes classification probabilities:
-where is the output representation of token from RoBERTa Zhuang et al. (2021); Aly et al. (2021). The loss for this task is the Binary Cross Entropy loss:
-3.7.2 Rationale Regularization Loss
-To encourage sparsity and continuity in rationale predictions, we introduce a Rationale Regularization Loss , which consists of:
-1. Sparsity Loss: Encourages the model to minimize the number of tokens selected as rationales:
-where is a mask ensuring that the loss is computed only on non-padding tokens.
-2. Continuity Loss: Encourages smoother rationale selection by minimizing sudden jumps between selected tokens:
-The total rationale regularization loss is:
-3.7.3 Combined Loss Function
-The final objective function integrates all components:
-where controls the weight of the rationale tagging loss, and , are hyperparameters for regularization.
+We adopt a two-stage classification framework to enhance claim verification robustness and mitigate label imbalance, especially the overrepresentation of NEI.
+Stage 1: Three-Class Classification.
+Given a claim-evidence pair , a BERT-based model predicts a probability distribution over three labels: Supported, Refuted, and Not Enough Information (NEI):
+This step is optimized using Cross-Entropy Loss.
+Stage 2: Binary Classification.
+If , we apply a refined binary classifier to distinguish between Supported and Refuted:
+This model uses Focal Loss Lin et al. (2018) to address class imbalance.
+Final Prediction Rule.
+The final verdict where is determined by comparing the confidence scores from both classifiers.
+Here, represents the index of the predicted label, where each index corresponds to a specific class description.
+where , represents the highest probability.
+This hybrid strategy allows the three-class model to handle general cases, especially detecting NEI early, while the binary model specializes in distinguishing difficult SUP/REF cases.
+3.4 SemViQA Pipeline System
+We now describe the full SemViQA pipeline, as illustrated in Figure 2.
+First, we prepare input for TF-IDF by splitting the context paragraph into sentences, then concatenating each sentence with the claim .
+We calculate the matching score for each sentence and select the one with the highest probability.
+If this score exceeds the threshold , we directly use this sentence as the evidence.
+If the score is below , we proceed to prepare input for the QATC model.
+We segment into subcontexts (as detailed in Section 3.1.1).
+Each subcontext is sequentially processed by the QATC model.
+If QATC identifies zero or multiple candidate evidence spans, we collect all predicted spans and re-rank them using TF-IDF.
+If QATC finds exactly one candidate evidence span, we confidently select it as the final evidence.
+Finally, we move to the two-step verdict classification (TVC) stage.
+We prepare input for TVC by concatenating the claim with the final evidence.
+We first apply the three-class model.
+If it predicts NEI, the process ends.
+Otherwise, we use an ensemble method to combine the weights from both the three-class and binary models to make the final prediction.
+All coefficients used in our system are provided in the ablation study, as shown in Appendix C.
 4 Experiments
 4.1 Dataset
-ISE-DSC01, used in the UIT Challenge 2023, is designed for AI evaluation in information extraction and verification.
-It comprises training, public test, and private test sets with diverse claim-context pairs, posing challenges in evidence retrieval and fact-checking.
-ViWikiFC Le et al.
-(2024) was developed to combat misinformation, featuring over 20,000 claims from Wikipedia.
-Unlike ISE-DSC01, it explicitly includes evidence for the “nei” (Not Enough Information) label, enhancing real-world applicability.
-| ISE-DSC01 | ViWikiFC | |
-| Train | 37,967 | 16,738 |
-| Dev | 4,794 | 2,090 |
-| Test | 5,396 | 2,091 |
-Table 1 summarizes the datasets. For ViWikiFC, experiments were conducted on all three sets: training, public test, and private test.
-| Method | Parameter | ViWikiFC | ISE-DSC01 | Avg Strict Acc | |||||||
-| ER | VC | Strict Acc | VC Acc | ER Acc | Time (s) | Strict Acc | VC Acc | ER Acc | Time (s) | ||
-| TF-IDF | InfoXLMlarge | 560M | 75.56 | 82.21 | 90.15 | 131 | 73.59 | 78.08 | 76.61 | 378 | 74.58 |
-| XLM-Rlarge | 560M | 76.47 | 82.78 | 90.15 | 134 | 75.61 | 80.50 | 78.58 | 366 | 76.04 | |
-| Ernie-Mlarge | 560M | 75.56 | 81.83 | 90.15 | 144 | 78.19 | 81.69 | 80.65 | 403 | 76.88 | |
-| BM25 | InfoXLMlarge | 560M | 70.44 | 79.01 | 83.50 | 130 | 72.09 | 77.37 | 75.04 | 320 | 71.27 |
-| XLM-Rlarge | 560M | 70.97 | 78.91 | 83.50 | 132 | 73.94 | 79.37 | 76.95 | 333 | 72.46 | |
-| Ernie-Mlarge | 560M | 70.21 | 78.29 | 83.50 | 141 | 76.58 | 80.76 | 79.02 | 381 | 73.40 | |
-| SBert | InfoXLMlarge | 838M | 74.99 | 81.59 | 89.72 | 195 | 71.20 | 76.59 | 74.15 | 915 | 73.10 |
-| XLM-Rlarge | 838M | 75.80 | 82.35 | 89.72 | 194 | 72.85 | 78.78 | 75.89 | 835 | 74.33 | |
-| Ernie-Mlarge | 838M | 75.13 | 81.44 | 89.72 | 203 | 75.46 | 79.89 | 77.91 | 920 | 75.30 | |
-| QA-based approaches | VC | ||||||||||
-| ViMRClarge | InfoXLMlarge | 1120M | 77.28 | 81.97 | 92.49 | 3778 | 54.36 | 64.14 | 56.84 | 9798 | 65.82 |
-| XLM-Rlarge | 1120M | 78.29 | 82.83 | 92.49 | 3824 | 53.98 | 66.70 | 57.77 | 9809 | 66.14 | |
-| Ernie-Mlarge | 1120M | 77.38 | 81.92 | 92.49 | 3785 | 56.62 | 62.19 | 58.91 | 9833 | 67.00 | |
-| InfoXLMlarge | InfoXLMlarge | 1120M | 78.14 | 82.07 | 93.45 | 4092 | 53.50 | 63.83 | 56.17 | 10057 | 65.82 |
-| XLM-Rlarge | 1120M | 79.20 | 83.07 | 93.45 | 4096 | 53.32 | 66.70 | 57.25 | 10066 | 66.26 | |
-| Ernie-Mlarge | 1120M | 78.24 | 82.21 | 93.45 | 4102 | 56.34 | 62.36 | 58.69 | 10078 | 67.29 | |
-| LLM | |||||||||||
-| Qwen2.5-1.5B-Instruct | 1.5B | 51.03 | 65.18 | 78.96 | 7665 | 59.23 | 66.68 | 65.51 | 19780 | 55.13 | |
-| Qwen2.5-3B-Instruct | 3B | 44.38 | 62.31 | 71.35 | 12123 | 60.87 | 66.92 | 66.10 | 31284 | 52.63 | |
-| LLM | VC | ||||||||||
-| Qwen2.5-1.5B-Instruct | InfoXLMlarge | 2B | 66.14 | 76.47 | 78.96 | 7788 | 64.40 | 68.37 | 66.49 | 19970 | 65.27 |
-| XLM-Rlarge | 2B | 67.67 | 78.10 | 78.96 | 7789 | 64.66 | 69.63 | 66.72 | 19976 | 66.17 | |
-| Ernie-Mlarge | 2B | 66.52 | 76.52 | 78.96 | 7794 | 65.70 | 68.37 | 67.33 | 20003 | 66.11 | |
-| Qwen2.5-3B-Instruct | InfoXLMlarge | 3.5B | 59.88 | 72.50 | 71.35 | 12246 | 65.72 | 69.66 | 67.51 | 31477 | 62.80 |
-| XLM-Rlarge | 3.5B | 60.74 | 73.08 | 71.35 | 12246 | 66.12 | 70.44 | 67.83 | 31483 | 63.43 | |
-| Ernie-Mlarge | 3.5B | 60.02 | 72.21 | 71.35 | 12251 | 67.48 | 70.77 | 68.75 | 31512 | 63.80 | |
-| SER Faster (ours) | TVC (ours) | ||||||||||
-| TF-IDF + ViMRClarge | Ernie-Mlarge | 1680M | 79.44 | 82.93 | 94.60 | 410 | 78.32 | 81.91 | 80.26 | 995 | 78.88 |
-| TF-IDF + InfoXLMlarge | 1680M | 79.77 | 83.07 | 95.03 | 487 | 78.37 | 81.91 | 80.32 | 925 | 79.07 | |
-| SER (ours) | TVC (ours) | ||||||||||
-| TF-IDF + ViMRClarge | InfoXLMlarge | 1680M | 80.25 | 83.84 | 94.69 | 2731 | 75.13 | 79.54 | 76.87 | 5191 | 77.69 |
-| XLM-Rlarge | 1680M | 80.34 | 83.64 | 94.69 | 2733 | 76.71 | 81.65 | 78.91 | 5219 | 78.53 | |
-| Ernie-Mlarge | 1680M | 79.53 | 82.97 | 94.69 | 2733 | 78.97 | 82.54 | 80.91 | 5225 | 79.25 | |
-| TF-IDF + InfoXLMlarge | InfoXLMlarge | 1680M | 80.68 | 83.98 | 95.31 | 3860 | 75.13 | 79.60 | 76.87 | 5175 | 77.91 |
-| XLM-Rlarge | 1680M | 80.82 | 83.88 | 95.31 | 3843 | 76.74 | 81.71 | 78.95 | 5200 | 78.78 | |
-| Ernie-Mlarge | 1680M | 80.06 | 83.17 | 95.31 | 3891 | 78.97 | 82.49 | 80.91 | 5297 | 79.52 | |
+We use two Vietnamese fact verification datasets: ISE-DSC01 from the UIT Challenge 2023 and ViWikiFC Le et al.
+(2024), which contains over 20,000 Wikipedia-based claims, including annotated evidence for the “nei” label.
+Dataset details are provided in Appendix A.1.
+| Method | ViWikiFC | ISE-DSC01 | Avg Strict Acc | |||||||
+| ER | VC | Strict Acc | VC Acc | ER Acc | Time (s) | Strict Acc | VC Acc | ER Acc | Time (s) | |
+| Traditional Baselines | ||||||||||
+| TF-IDF | InfoXLM | 75.56 | 82.21 | 90.15 | 131 | 73.59 | 78.08 | 76.61 | 378 | 74.58 |
+| XLM-R | 76.47 | 82.78 | 90.15 | 134 | 75.61 | 80.50 | 78.58 | 366 | 76.04 | |
+| Ernie-M | 75.56 | 81.83 | 90.15 | 144 | 78.19 | 81.69 | 80.65 | 403 | 76.88 | |
+| BM25 | InfoXLM | 70.44 | 79.01 | 83.50 | 130 | 72.09 | 77.37 | 75.04 | 320 | 71.27 |
+| XLM-R | 70.97 | 78.91 | 83.50 | 132 | 73.94 | 79.37 | 76.95 | 333 | 72.46 | |
+| Ernie-M | 70.21 | 78.29 | 83.50 | 141 | 76.58 | 80.76 | 79.02 | 381 | 73.40 | |
+| SBert | InfoXLM | 74.99 | 81.59 | 89.72 | 195 | 71.20 | 76.59 | 74.15 | 915 | 73.10 |
+| XLM-R | 75.80 | 82.35 | 89.72 | 194 | 72.85 | 78.78 | 75.89 | 835 | 74.33 | |
+| Ernie-M | 75.13 | 81.44 | 89.72 | 203 | 75.46 | 79.89 | 77.91 | 920 | 75.30 | |
+| QA-based Approaches | ||||||||||
+| ViMRC | InfoXLM | 77.28 | 81.97 | 92.49 | 3778 | 54.36 | 64.14 | 56.84 | 9798 | 65.82 |
+| XLM-R | 78.29 | 82.83 | 92.49 | 3824 | 53.98 | 66.70 | 57.77 | 9809 | 66.14 | |
+| Ernie-M | 77.38 | 81.92 | 92.49 | 3785 | 56.62 | 62.19 | 58.91 | 9833 | 67.00 | |
+| InfoXLM | InfoXLM | 78.14 | 82.07 | 93.45 | 4092 | 53.50 | 63.83 | 56.17 | 10057 | 65.82 |
+| XLM-R | 79.20 | 83.07 | 93.45 | 4096 | 53.32 | 66.70 | 57.25 | 10066 | 66.26 | |
+| Ernie-M | 78.24 | 82.21 | 93.45 | 4102 | 56.34 | 62.36 | 58.69 | 10078 | 67.29 | |
+| LLMs | ||||||||||
+| Qwen2.5-1.5-Instruct | 51.03 | 65.18 | 78.96 | 7665 | 59.23 | 66.68 | 65.51 | 19780 | 55.13 | |
+| Qwen2.5-3B-Instruct | 44.38 | 62.31 | 71.35 | 12123 | 60.87 | 66.92 | 66.10 | 31284 | 52.63 | |
+| Qwen2.5-1.5-Instruct | InfoXLM | 66.14 | 76.47 | 78.96 | 7788 | 64.40 | 68.37 | 66.49 | 19970 | 65.27 |
+| XLM-R | 67.67 | 78.10 | 78.96 | 7789 | 64.66 | 69.63 | 66.72 | 19976 | 66.17 | |
+| Ernie-M | 66.52 | 76.52 | 78.96 | 7794 | 65.70 | 68.37 | 67.33 | 20003 | 66.11 | |
+| Qwen2.5-3B-Instruct | InfoXLM | 59.88 | 72.50 | 71.35 | 12246 | 65.72 | 69.66 | 67.51 | 31477 | 62.80 |
+| XLM-R | 60.74 | 73.08 | 71.35 | 12246 | 66.12 | 70.44 | 67.83 | 31483 | 63.43 | |
+| Ernie-M | 60.02 | 72.21 | 71.35 | 12251 | 67.48 | 70.77 | 68.75 | 31512 | 63.80 | |
+| Ours: SER Faster + TVC | ||||||||||
+| ViMRC | Ernie-M | 79.44 | 82.93 | 94.60 | 410 | 78.32 | 81.91 | 80.26 | 995 | 78.88 |
+| InfoXLM | 79.77 | 83.07 | 95.03 | 487 | 78.37 | 81.91 | 80.32 | 925 | 79.07 | |
+| Ours: Full SER + TVC | ||||||||||
+| ViMRC | InfoXLM | 80.25 | 83.84 | 94.69 | 2731 | 75.13 | 79.54 | 76.87 | 5191 | 77.69 |
+| XLM-R | 80.34 | 83.64 | 94.69 | 2733 | 76.71 | 81.65 | 78.91 | 5219 | 78.53 | |
+| Ernie-M | 79.53 | 82.97 | 94.69 | 2733 | 78.97 | 82.54 | 80.91 | 5225 | 79.25 | |
+| InfoXLM | InfoXLM | 80.68 | 83.98 | 95.31 | 3860 | 75.13 | 79.60 | 76.87 | 5175 | 77.91 |
+| XLM-R | 80.82 | 83.88 | 95.31 | 3843 | 76.74 | 81.71 | 78.95 | 5200 | 78.78 | |
+| Ernie-M | 80.06 | 83.17 | 95.31 | 3891 | 78.97 | 82.49 | 80.91 | 5297 | 79.52 | |
 4.2 Experimental Setup
-We optimized our model through extensive experiments on an NVIDIA A100 GPU, fine-tuning key hyperparameters while maintaining consistent settings across all runs.
-The final hyperparameters, selected via rigorous validation, improved both accuracy and strict accuracy on ISE-DSC01 and ViWikiFC.
-Details are provided in Appendix B.
-For evaluation, all methods were tested on an NVIDIA T4 instance in Kaggle to ensure fairness.
-Additionally, the large language model was fine-tuned on a distributed setup with A100 GPUs, utilizing a structured prompt reformulation process.
-The dataset, originally in raw text and labels, was transformed into a prompt-based format optimized for LLM fine-tuning.
-This ensured effective task-specific learning while aligning with pre-trained knowledge.
-Further details on training setup, prompt design, and preprocessing are in Appendix B.
+We conducted extensive experiments on NVIDIA A100 GPUs, fine-tuning key hyperparameters while keeping consistent settings across runs.
+The final configuration, selected via rigorous validation, improved both accuracy and strict accuracy on ISE-DSC01 and ViWikiFC.
+Full details are provided in Appendix D.
+For fair evaluation, all methods were tested on a Kaggle instance with an NVIDIA T4 GPU.
+The large language model was fine-tuned in a distributed A100 setup using a structured prompt-based reformulation.
+Raw data were converted into prompt format to align with LLM training objectives and maximize task-specific performance.
+Training setup, prompt design, and preprocessing are also detailed in Appendix D.
 4.3 Main Results
@@
 a) Handling Long Token Sequences in Fact-Checking
-One major limitation of conventional Question Answering (QA) models in fact verification tasks is their inability to process long-context claims due to token length constraints.
-Transformer-based models such as ViMRClarge, InfoXLMlarge Chi et al.
-(2021), and XLM-Rlarge Conneau et al.
-(2020), Ernie-Mlarge Ouyang et al.
-(2021) are restricted to a maximum input length of 512 tokens.
-However, real-world fact-checking datasets, such as ISE-DSC01, often contain contexts exceeding 4800 tokens, causing a significant drop in performance for QA-based approaches due to their inability to handle long-token sequences.
-This limitation prevents these models from effectively leveraging the full evidence available, leading to incomplete or inaccurate fact verification.
-To address this issue, our proposed SemViQA model incorporates a retrieval-based approach capable of handling long-token sequences efficiently, as illustrated in Figure 3.
-On ISE-DSC01, where long contexts are essential for verification, SemViQA significantly outperforms traditional QA models by effectively capturing and utilizing extended contextual information.
-This demonstrates that the long-token issue is a critical bottleneck in fact verification tasks.
-In contrast, on the ViWikiFC dataset, where the token length is closer to 512 tokens, QA-based models achieve competitive performance, further confirming that the long-token issue is a major factor in performance degradation.
-However, when the long-token limitation does not exist on the ViWikiFC dataset, integrating our SER approach into a QA-based model proves to be an effective solution.
-By leveraging SER, we achieve approximately 1.86% higher ER Accuracy on ViWikiFC, demonstrating its efficiency in optimizing fact verification when token constraints are minimized.
-These results highlight that long-token constraints are a fundamental challenge in fact verification.
-Our approach not only mitigates this issue through retrieval-based processing in SemViQA but also enhances QA models with QACT when long-token constraints are less severe, ensuring improved performance across diverse datasets.
+A major limitation of conventional Question Answering (QA) models in fact verification is their inability to process long-context claims due to the 512-token input limit of transformer-based models such as ViMRClarge333https://huggingface.co/nguyenvulebinh/vi-mrc-large, InfoXLMlarge Chi et al.
+(2021), XLM-Rlarge Conneau et al.
+(2020), and Ernie-Mlarge Ouyang et al.
+(2021).
+Real-world datasets like ISE-DSC01 often contain contexts exceeding 4800 tokens, severely degrading QA-based performance by limiting access to full evidence.
+To overcome this, SemViQA employs an efficient retrieval-based strategy (see Section 3.1.1) that handles long-token sequences effectively.
+On ISE-DSC01, SemViQA outperforms traditional QA models by fully leveraging extended contexts, confirming that the long-token constraint is a critical bottleneck.
+Conversely, on ViWikiFC, where contexts average around 512 tokens, QA models perform competitively.
+Yet, even in this setting, integrating our Semantic-based Evidence Retrieval (SER) yields a 1.86% improvement in evidence retrieval accuracy, demonstrating the versatility and efficiency of our approach.
+These findings emphasize that long-token limitations significantly hinder fact verification, and SemViQA successfully mitigates this issue while enhancing QA models across varying dataset conditions.
 b) Performance and Inference Time Optimization
-One of the key advancements of SemViQA over other methods is its significant reduction in inference time without compromising accuracy. Key observations include:
-The average inference time of SemViQA on ISE-DSC01 is 5200s, whereas large LLM-based models such as Qwen2.5-3B-Instruct Qwen et al.
-(2024) require over 31,000s, making SemViQA at least 6 times faster than these large language model approaches.
-Compared to ViMRClarge333https://huggingface.co/nguyenvulebinh/vi-mrc-large (9800s on ISE-DSC01), SemViQA reduces inference time by nearly 50% while maintaining higher performance in both Strict Accuracy and Veracity Classification Accuracy.
-Methods based on SBERT Reimers and Gurevych (2019b) or BM25 have lower inference times but fail to maintain high accuracy, particularly when dealing with multi-step reasoning tasks.
-SemViQA achieves a better balance between classification performance and processing speed, ensuring practical deployment in real-world environments.
-SemViQA Faster: We introduce an optimized version of SemViQA that achieves up to 7× faster inference while maintaining high accuracy.
-As shown in Figure 4, the inference speed of SER-Faster closely matches that of traditional methods while still outperforming competing solutions in both efficiency and accuracy.
-To achieve this speed improvement, we process all subcontexts within a sample as a batch, enabling batch inference (see Section 3.1 for details on subcontexts).
-While this approach results in a slight performance drop compared to the standard SemViQA model, the trade-off is minimal considering the significant reduction in processing time.
-This efficiency enhancement makes SemViQA Faster highly suitable for real-world deployment, allowing seamless integration into large-scale fact-checking applications.
-4.3.2 Analysis of Confidence Threshold in SemViQA
-The confidence threshold plays a crucial role in balancing accuracy and inference time in SemViQA’s evidence retrieval process.
-Analysis from Figure 5 indicates that as the threshold increases from 0.0 to 0.5, evidence retrieval accuracy improves significantly, particularly on ViWikiFC ( 95%) and ISE-DSC01 ( 80.8%).
-However, beyond 0.5, accuracy gains plateau, while inference time decreases sharply due to the system filtering out low-confidence evidence more aggressively.
-Setting an optimal threshold in the range of 0.4 – 0.5 achieves a trade-off between efficiency and accuracy, ensuring that SemViQA operates swiftly while maintaining precise evidence retrieval.
+SemViQA significantly reduces inference time while maintaining high accuracy, making it highly practical for real-world applications. Key highlights include:
+On ISE-DSC01, SemViQA averages 5200s per run, over 6 times faster than large LLM-based models like Qwen2.5-3B-Instruct Qwen et al. (2024), which require over 31,000s.
+Compared to ViMRClarge (9800s), SemViQA halves inference time while achieving superior Strict Accuracy and Veracity Classification Accuracy.
+Although BM25 and SBERT Reimers and Gurevych (2019b) are faster, they struggle with complex, multi-step reasoning where SemViQA maintains a strong balance between speed and accuracy.
+SemViQA Faster: We further introduce SemViQA Faster, which accelerates inference by up to 7 through batch processing of subcontexts (see Section 3.1.1).
+As shown in Figure 4, this variant achieves inference speeds comparable to traditional methods while retaining competitive accuracy.
+The minor performance trade-off is acceptable given the substantial time savings, making SemViQA Faster ideal for scalable, real-world fact-checking systems.
 4.3.2 Comparison with other results in the competition
@@
 | SemViQA | 78.97 | 82.54 | 80.91 |
-| DS@UIT Dynasty | 78.05 | 84.76 | 80.13 |
-| URA_FNU | 77.87 | 83.71 | 79.96 |
-| Plain Sailing | 77.09 | 82.33 | 78.31 |
-| ViNSV | 76.33 | 81.67 | 78.11 |
+| DS@UIT Dynasty444https://github.com/minhquan6203/Verdict-Classification-for-Fact-Checking-at-DSC-2023 | 78.05 | 84.76 | 80.13 |
+| URA_FNU555https://github.com/virrosluo/URA_UIT_Data_Science_Challenge | 77.87 | 83.71 | 79.96 |
+| ViNSV Tran et al. (2024b) | 76.33 | 81.67 | 78.11 |
+| Tran et al. (2024a) | 75.11 | 82.30 | 76.82 |
 The results presented in Table 2 indicate that our SemViQA approach outperforms other competing teams, achieving the highest Strict Accuracy and demonstrating exceptional effectiveness in information processing and verification.
@@
 Future work should focus on refining retrieval mechanisms and classification strategies to enhance efficiency and robustness, ensuring broader applicability of SemViQA in real-world fact verification scenarios.
-Appendix A Strict Accuracy in Fact-Checking
+Appendix A Comprehensive Data Processing
+A.1 Dataset Statistics
+This section presents the detailed statistics of the two datasets used in our experiments: ISE-DSC01 and ViWikiFC.
+The table below summarizes the number of samples in the training, development, and test sets for both corpora.
+These statistics help contextualize the scale and evaluation coverage of our methods.
+| ISE-DSC01 | ViWikiFC | |
+| Train | 37,967 | 16,738 |
+| Dev | 4,794 | 2,090 |
+| Test | 5,396 | 2,091 |
+A.2 Data Analysis on Context Lengths
+Figure 5 illustrates the distribution of token lengths across input contexts in the ViWikiFC and ISE-DSC01 datasets.
+As shown, many samples significantly exceed the 512-token input limitation of standard Transformer models.
+The ISE-DSC01 dataset, in particular, contains several contexts with over 4,000 tokens.
+This analysis highlights the necessity for effective context segmentation strategies to ensure full coverage of relevant evidence while maintaining compatibility with model constraints.
+Figure 3 illustrates our strategy for handling long input contexts.
+First, the context is segmented into individual sentences.
+Next, sentences are sequentially aggregated into subcontexts until reaching approximately 400 tokens.
+Each completed subcontext is then passed to the QATC model to identify potential evidence.
+The next subcontext begins from the subsequent sentence, continuing until all sentences have been processed.
+However, processing subcontexts sequentially can be time-consuming.
+Therefore, we developed SER Faster, which batches subcontexts and processes them in parallel, significantly accelerating the retrieval process.
+Appendix B Strict Accuracy in Fact-Checking
 Strict Accuracy: This metric is a stringent measure that requires both the verdict and the evidence to be predicted correctly compared to the ground truth sample.
@@
 Strict accuracy is the average of all StrAcc values.
-Appendix B Hyperparameter and LLM Training Configuration
+Appendix C Ablation Study
+C.1 Ablation Experiments
+| Configuration | ViWikiFC | ISE-DSC01 | |||||
+| ER Model | VC Model | Strict Acc | VC Acc | ER Acc | Strict Acc | VC Acc | ER Acc |
+| SemViQA (Full) | |||||||
+| InfoXLM | InfoXLM | 80.68 | 83.98 | 95.31 | 75.13 | 79.60 | 76.87 |
+| XLM-R | 80.82 | 83.88 | 95.31 | 76.74 | 81.71 | 78.95 | |
+| Ernie-M | 80.06 | 83.17 | 95.31 | 78.97 | 82.49 | 80.91 | |
+| w/o Binary Classification (One-step VC) | |||||||
+| InfoXLM | InfoXLM | 79.63 | 82.88 | 95.31 | 73.87 | 78.35 | 76.89 |
+| XLM-R | 80.73 | 83.69 | 95.31 | 75.96 | 80.80 | 78.97 | |
+| Ernie-M | 79.91 | 83.07 | 95.31 | 78.47 | 81.89 | 80.93 | |
+| w/o QATC (TF-IDF-based ER) | |||||||
+| TF-IDF | InfoXLM | 76.57 | 83.26 | 90.15 | 74.89 | 79.36 | 76.61 |
+| XLM-R | 76.47 | 82.93 | 90.15 | 76.39 | 81.41 | 78.58 | |
+| Ernie-M | 75.75 | 81.97 | 90.15 | 78.71 | 82.28 | 80.65 | |
+Table 4 presents the ablation results to evaluate the contribution of each component in the SemViQA framework.
+When employing the full model with both the QATC-based evidence retrieval and the two-step verdict classification (TVC), SemViQA achieves the best performance across both datasets.
+Notably, using Ernie-Mlarge yields the highest strict accuracy (78.97%) and evidence retrieval accuracy (80.91%) on the ISE-DSC01 dataset.
+Removing the binary classification stage (i.e., using only one-step VC) leads to a noticeable performance drop, especially on ISE-DSC01, indicating that the binary classifier enhances the distinction between Supported and Refuted labels.
+Furthermore, replacing the QATC module with a TF-IDF based retriever results in a significant decline (about 5%) in evidence retrieval accuracy, which subsequently affects the overall performance.
+These findings highlight the critical role of both QATC and the two-step classification scheme in improving SemViQA’s effectiveness on fact verification in Vietnamese.
+C.2 Analysis of Confidence Threshold in SemViQA
+The confidence threshold plays a crucial role in balancing accuracy and inference time in SemViQA’s evidence retrieval process.
+Analysis from Figure 6 indicates that as the threshold increases from 0.0 to 0.5, evidence retrieval accuracy improves significantly, particularly on ViWikiFC ( 95%) and ISE-DSC01 ( 80.8%).
+However, beyond 0.5, accuracy gains plateau, while inference time decreases sharply due to the system filtering out low-confidence evidence more aggressively.
+Setting an optimal threshold in the range of 0.4 - 0.5 achieves a trade-off between efficiency and accuracy, ensuring that SemViQA operates swiftly while maintaining precise evidence retrieval.
+C.3 Effectiveness of QATC over Traditional QA-based Evidence Retrieval
+| SemViQA | ViWikiFC | ISE-DSC01 | |||||
+| SER | TVC | Strict Acc | VC Acc | ER Acc | Strict Acc | VC Acc | ER Acc |
+| QATC-based ER | |||||||
+| InfoXLM | InfoXLM | 80.68 | 83.98 | 95.31 | 75.13 | 79.60 | 76.87 |
+| XLM-R | 80.82 | 83.88 | 95.31 | 76.74 | 81.71 | 78.95 | |
+| Ernie-M | 80.06 | 83.17 | 95.31 | 78.97 | 82.49 | 80.91 | |
+| ViMRClarge | InfoXLMlarge | 80.25 | 83.84 | 94.69 | 75.13 | 79.54 | 76.87 |
+| XLM-Rlarge | 80.34 | 83.64 | 94.69 | 76.71 | 81.65 | 78.91 | |
+| Ernie-Mlarge | 79.53 | 82.97 | 94.69 | 78.97 | 82.54 | 80.91 | |
+| QA-based ER | |||||||
+| InfoXLM | InfoXLM | 79.96 | 83.50 | 94.45 | 74.02 | 78.95 | 75.83 |
+| XLM-R | 80.11 | 83.60 | 94.45 | 75.61 | 80.95 | 77.91 | |
+| Ernie-M | 79.24 | 82.74 | 94.45 | 77.82 | 81.76 | 79.82 | |
+| ViMRClarge | InfoXLMlarge | 79.77 | 83.84 | 94.26 | 74.05 | 78.93 | 75.87 |
+| XLM-R | 79.87 | 83.79 | 94.26 | 75.65 | 80.93 | 77.95 | |
+| Ernie-M | 79.01 | 82.78 | 94.26 | 77.84 | 81.73 | 79.86 | |
+To compare the learning capabilities of QATC with traditional QA models, we construct two versions of the SemViQA pipeline that are identical in structure, differing only in the evidence retrieval model.
+As shown in Table 5, using QATC consistently outperforms QA-based retrieval across both datasets.
+Specifically, on ViWikiFC, QATC achieves up to 80.82% Strict Accuracy and 95.31% ER Accuracy, while QA-based models peak at 80.11% and 94.45%, respectively.
+The improvement is even more evident on ISE-DSC01, where QATC reaches 78.97% Strict Accuracy and 80.91% ER Accuracy.
+These results confirm that QATC is more effective at learning to identify relevant evidence, leading to superior performance across the entire fact-checking pipeline.
+C.4 Analysis of Confusion Matrix in Verdict Classification
+To evaluate the effectiveness of our Two-step Verdict Classification (TVC) strategy, we analyze the confusion matrices in Figure 7, which compare the standard three-class classification with our proposed two-step approach, focusing solely on Supported and Refuted claims.
+The three-class classifier achieves strong performance, with accuracy of 95.4% for Supported and 94.6% for Refuted; however, it still exhibits notable confusion between the two classes, likely due to their semantic proximity and shared evidence patterns.
+In contrast, the two-step approach—by isolating NEI cases early and applying a dedicated binary classifier further improves accuracy to 98.0% for Supported and 96.4% for Refuted.
+These results support our hypothesis that decomposing the verdict classification task into semantically coherent subtasks enhances the model’s precision in detecting factual consistency.
+Appendix D Hyperparameter and LLM Training Configuration
 In this section, we present the detailed hyperparameter settings and training configurations for both our SemViQA models and the Large Language Model (LLM) fine-tuning process.
-Table 4 consolidates all hyperparameters used across different models, including Binary Classification (BC), Three-Class Classification (TC), Question Answering Token Classifier (QATC), and LLM fine-tuning.
+Table 6 consolidates all hyperparameters used across different models, including Binary Classification (BC), Three-Class Classification (TC), Question Answering with Token Classification (QATC), and LLM fine-tuning.
 | Hyperparameter | BC | TC | QATC | LLM |
@@
 These datasets have been carefully adapted for training to improve performance and ensure compatibility with our model.
-For training, we utilize the official Qwen LLM implementation from the QwenLM repository444https://github.com/QwenLM/Qwen.
+For training, we utilize the official Qwen LLM implementation from the QwenLM repository666https://github.com/QwenLM/Qwen.
 Our training setup follows the full configuration outlined in Table 6, ensuring optimal efficiency and alignment with best practices.
-[backgroundcolor=gray!5, linewidth=0.75pt, roundcorner=5pt]
-Question: You are tasked with verifying the correctness of the following statement.
 - We provide you with a claim and a context. Please classify the claim into one of three labels: “Supported”, “REFUTED”, or “NEI” (Not Enough Info).
@@
 - Remember, the evidence must be a full sentence, not part of a sentence or less than one sentence.
-Given a claim and context as follows:
-Context: The actress revealed her secrets to maintaining a youthful appearance as follows: Eating three balanced meals a day.
+Given a claim and context as follows: Context: The actress revealed her secrets to maintaining a youthful appearance as follows: Eating three balanced meals a day.
 For dinner, Ivy Chen usually eats early to ensure her body has enough time to digest food, metabolize energy, and avoid putting pressure on the stomach and other organs.
@@
 Claim: Even when traveling abroad, Ivy Chen maintains her running habit.
-Answer: This claim is classified as SUPPORTED. The evidence is: Even when traveling abroad, she maintains her running habit.
+Answer: This claim is classified as Supported.
+The evidence is: Even when traveling abroad, she maintains her running habit.
 We present the complete training progress of the LLM models and QATC in Figure 9 and Figure 8, respectively.
@@
 These results highlight the convergence behavior of QATC training across different architectures, further supporting the robustness of our approach.
-Appendix C Comparison of TF-IDF and QATC in Fact-Checking: Examples of Incorrect vs. Correct Evidence Selection
+Appendix E Comparison of TF-IDF and QATC in Fact-Checking: Examples of Incorrect vs. Correct Evidence Selection
 | Claim | Evidence | TF-IDF | QATC |